# Visual Highlighting for Situated Brushing and Linking

### v1

@@ 1 Introduction
Situated analytics (SitA) involves the use of situated visualizations [WF09] to improve sense-making in physical contexts [$*{}^{*}$15, $*{}^{*}$18]. It is facilitated predominantly by the use of augmented reality (AR) [$*{}^{*}$22]. Willett et al. [WJD17] distinguish visualizations directly embedded with one particular physical referent from situated visualizations that indirectly link to a referent or the entire environment. Shin et al. [$*{}^{*}$23] note that most current SitA tools use the referent as a trigger to instantiate and modify visualizations. For instance, scanner applications augment real-world objects with additional information as they come into view. SitA is therefore reliant on the user’s viewpoint and interactions with the real world to determine what information to display.
Although this “physical world first” approach feels natural in the context of AR, it is inherently constrained in situations with many physical referents or referents that are spread far apart [LSS23]. Consider the classic supermarket scenario introduced by ElSayed et al. [$*{}^{*}$16], which features each product in the store having its own embedded visualization. While this scenario works well for products in view, it does not scale well to products outside of one’s view. Even with out-of-view labeling techniques [$*{}^{*}$18, LYBP23], the customer in one-to-one configuration would still need to look at and interpret the visualization of each individual product [LSS23]. The issue is exacerbated by the small field of view (FOV) of current state-of-the-art AR headsets.
@@ 2.1 Immersive and situated analytics
Immersive technologies, both VR and AR, are increasingly being used for immersive analytics [$*{}^{*}$18] of spatial data [FP21] and abstract data [$*{}^{*}$22]. Research has shown how the 3D environment can be used as an “immersive space to think” [$*{}^{*}$21], populating this space with multiple views and representations [RBR22]. However, care must be taken to consider the pitfalls of designing for immersive displays [ML14]. Designers of VR interfaces for analytics must take into account the characteristics of VR displays that differ from desktop computing, such as stereoscopic depth, angular resolution, field of view, or fatigue related to physical navigation and interaction [$*{}^{*}$17].
Recent work considers immersive analytics techniques with AR in physical environments, a new direction referred to as situated analytics (SitA) [$*{}^{*}$15, $*{}^{*}$16]. SitA is characterized by the ability to perform higher-level analytical reasoning aided by the user’s physical environment [$*{}^{*}$23], particularly when there is a semantic relationship between the data and its physical referents [WF09, WJD17]. While research has investigated how situated visualizations can be designed for SitA [LSS23, $*{}^{*}$23], considering human perception to leverage efficient visual processing—particularly in situated contexts—is a grand challenge [$*{}^{*}$21].
Designers of SitA systems (and AR interfaces in general) must handle the additional constraint that visual representations cannot always be positioned on a “clean slate” background. Kruijff et al. [KSF10] list several challenges related to visual perception in AR which are caused by the environment, image capture, display, and visual representation. Satkowski et al. [SD21] study how the real-world background affects perception in AR, and Assor et al. [APHD24] propose a design space of how to handle non-visible physical referents. Similarly, AR must always consider the interplay of visual representation with the existing and immutable appearance of the real world. Therefore, any visualization pipeline targeting AR needs to deal with the fusion of real and virtual visual attributes [$*{}^{*}$21]. Addressing these challenges requires a wide range of visual encoding methods [SH16, chapter 7].
@@ 2.2 Brushing and linking
One of the core topics of this paper is brushing and linking [Rob07], which is the process of locating the corresponding items of interest in multiple views. Fundamentally, it comprises two sequential stages, wherein selections made in one visualization (brushing) are automatically shown in another visualization (linking) [Kei02]. Much research has focused on the brushing component, investigating interaction techniques to support the effective selection of data points in desktop computing [$*{}^{*}$18].
In immersive analytics, brushing and linking have seen (arguably limited) use, be it to compare multiple immersive coordinated views [$*{}^{*}$15, LPED20] or to support collaborative awareness [$*{}^{*}$21, SBDE23, SS22]. Most methods use simple 3D selection techniques (e.g., raycasting) or pure 2D selection techniques (e.g., touchscreen inputs). While we do not propose new brushing interaction techniques, we are the first (to the best of our knowledge) to explore brushing and linking in a SitA context—particularly when using physical referents as a coordinated view.
@@ 2.3 Attention guidance in augmented and virtual reality
In SitA, we assume that the linking component targets views and referents in the real world. Hence, it can be described as a visual search task for target objects in an AR display. Note that this work considers attention guidance as a means to support visual search and identification, and not as instructional cues for movement training [YLS24] or assembly tasks [$*{}^{*}$23]. Wolfe et al. [WVEG11] characterize visual search as the task of visually identifying a target among other, irrelevant objects (distractors). If there are too many distractors, a naive visual search becomes inefficient [Ver02]. We must therefore ensure that the visual search is tolerant of clutter and occlusion, since we cannot change the real world. There should be sufficient contrast and legibility between real and virtual objects. We have to be able to direct a user’s attention to the relevant objects, in particular, if these are out of view.
@@
According to Cockburn et al. [CKB09], conventional InfoVis methods to guide visual attention and search, such as overview + detail or focus + context, occupy a lot of screen space as they enlarge objects or introduce replicas of different sizes. As a major alternative, they mention cue-based approaches, which modify object rendering styles. Lin et al. [LYBP23] note that such cues are preferred in AR, since resizing or duplicating occupies too much of the natural view space. These cue-based techniques for highlighting focal objects can select from a variety of rendering attributes. A baseline is formed by highlights, which emphasize a target by modifying its color throughout the target area or as an outline [$*{}^{*}$18].
@@
In immersive displays, several more sophisticated techniques have been proposed to ensure that a given contrast is achieved [GAM18, $*{}^{*}$13]. Contrast modulation, even below a detectable threshold [VMFS11], can pre-attentively guide the user. Other attributes that can be modified to highlight or attract attention include spatial frequencies (i.e., focus or blur) [KMH01], stereoscopy [KCWK20], or temporal frequencies [LSGB20].
@@
Cockburn et al. [CKB09] further mention proxies as cues. Arrow-shaped proxies are commonly used in AR [$*{}^{*}$18, TK06], most likely because they are simple to synthesize and easy to understand. Naturally, proxies are a popular means to guide toward out-of-view targets. In addition to arrows, various popular shapes include halos [BR03], funnels [BTOX06], compass-like [SL00] or radar-like widgets [JHPR11], and lines or trajectories [$*{}^{*}$19]. Previous work also compares the effect of various visual cues in VR and AR. Several authors [BSEN18, $*{}^{*}$14, $*{}^{*}$19, WGRF22] investigate the performance differences of visual search with respect to the field of view of the display. Other studies compare multiple attention guidance (i.e., highlighting) techniques with respect to the effectiveness and preservation of immersion [LSGB20], the effectiveness of multimodal cues [$*{}^{*}$20], or the influence of stationary and moving distractors [DARS23]. Moreover, our selection of techniques was influenced by the work of Whitlock et al. [WSS20] on the perception of fundamental graphical attributes in immersive analytics. However, none of these works investigate the use of AR or VR attention guidance in the context of brushing and linking. Therefore, our work seeks to bridge the gap between these two topics necessary to enable brushing and linking in SitA.
@@ 3.1 Scenario
SitA is relevant in many situations involving data and its physical context [$*{}^{*}$22]. Brushing and linking becomes useful when considering the actual appearance or location of physical referents. We therefore chose a supermarket scenario to situate our experimental prototype and the subsequent user study (Visual Highlighting for Situated Brushing and Linking). A supermarket is a popular scenario in the SitA literature [$*{}^{*}$15, $*{}^{*}$15, EST16]. Its defining feature is an abundance of grocery products (referents) laid out on shelves, which vary in layout, making the scenario well-suited for brushing and linking. We also considered similar environments, such as a library with books or a building façade with windows as referents [$*{}^{*}$16]. We built the supermarket using VR as a kind of information-rich virtual environment [$*{}^{*}$03, PKB05], serving as a simulated test environment that can be carefully controlled and replicated [RWBH09]. While using AR would be more ecologically valid, the use of VR avoids the technical challenges of tracking each product [CFSS23] or logistical challenges of renting or building a store. Previous work also suggested that findings in AR can be replicated in VR [GCOK21, LBHB09, LBBH10, $*{}^{*}$13], which further motivates and validates our use of a simulated supermarket in VR.
@@ 3.2 Highlighting techniques
Color is likely the de facto highlighting technique used in regular brushing and linking [$*{}^{*}$18, Rob07]. It mainly benefits from the pre-attentive “popout” effect, particularly when the new color highly contrasts with the background. Hence, we incorporate it into our prototype, changing the color of the product to a solid yellow when selected. While we had initially tried to use a semi-transparent yellow tint on the products instead, pilot tests showed that this mode was not able to sufficiently “pop-out” specific products: A red cereal box, when tinted yellow, simply looks like an orange cereal box, and not a red one that is highlighted.
@@
Outline is another common technique found in brushing and linking [$*{}^{*}$14, $*{}^{*}$14]. Moreover, it is also a popular technique used in visual highlighting to show that an object is “selected” [$*{}^{*}$18, $*{}^{*}$20]. Although Color and Outline have similar characteristics in how they highlight targets, the actual appearance of a product is visible for Outline but not for Color. Sometimes a more subtle, non-obtrusive highlighting can be beneficial (e.g., for safety). We show outlines on the borders of the selected grocery products in camera space. We chose a width for the outline which, after pilot testing, we felt was sufficiently wide to see the outline, but thin enough not to obstruct neighboring products.
@@
Links are trajectories used to draw connections between entities [CC07]. In the context of brushing and linking, a link indicates that two connected data marks are the same underlying data record [$*{}^{*}$18]. The same premise holds in SitA, where a data mark on the situated visualization (i.e., the virtual tablet in Visual Highlighting for Situated Brushing and Linking) is explicitly connected to its associated referent. Particularly for the comparison and analysis across view, direct connections can facilitate tasks. To this end, we leverage a prior Unity implementation by Prouzeau et al. [$*{}^{*}$19] which draws visual links between pairs of objects in a 3D scene. We adjusted some of their script’s parameters to make the links more responsive to movement, because we allow the situated visualization to be moved by the user.
@@
Arrows are ubiquitous symbols used to direct attention. They are popular for attention guidance in AR, particularly towards objects or points of interest that are out of view [KR14, TK06, WGRF22, $*{}^{*}$20]. Thus, we consider how similar arrows could be used to highlight objects in SitA. For our study, we extended FlyingARrow [$*{}^{*}$18], which repeatedly instantiates 3D arrows in front of the user that then fly toward a target object. Static arrows like Wieland et al. [WGRF22] or Yu et al. [$*{}^{*}$20] spawn in the viewport, causing clutter when there are multiple targets and hence might be less beneficial in real-life scenarios than dynamic arrows [$*{}^{*}$18]. We also suspect that adding animations or motion makes targets more salient. Since Gruenefeld et al. [$*{}^{*}$18] claim a low workload, we suspected the dynamic arrow to be a reasonable choice for highlighting in SitA. Our adapted version of FlyingARrow is able to present multiple arrows concurrently and dynamically change the arrows to support on-the-fly brushing.
@@ 3.3 Brushing interactions
As the premise of our work involves both brushing and linking, our prototype needs faculties to enable brushing in a situated visualization. A virtual tablet attached to the non-dominant-hand controller (Visual Highlighting for Situated Brushing and Linking, center) presents a 2D scatterplot, generated with DXR [$*{}^{*}$19]. The $x$ and $y$ data dimensions of the scatterplot are determined by buttons on the right side of the tablet. The tablet’s size (40 cm by 31 cm) ensures that its contents are easily readable in VR, while still small enough to not be visually obtrusive.
We opt for a simple brushing approach, reminiscent of the brushing mechanism used in FIESTA [$*{}^{*}$21]. Brushing is performed by pointing the dominant-hand controller at the scatterplot and pressing either the index button to add to the selection or the grip button to remove it from the selection. We provide two modes for brushing: spherical and rectangular. The mode can be changed using the corresponding buttons on the tablet. Any brushed data marks on the scatterplot turn a solid yellow. To make selections even easier, we added simple range filters along the two data axes. These filters are represented as handles which can be dragged using raycast interactions, similar to ImAxes [$*{}^{*}$17]. Filtered data marks cannot be brushed, ensuring that no unwanted points are included. Step-by-step images of this situated brushing and linking process are shown in Figure 1. The tablet could be held comfortably at hip height without obscuring any referents, as was the case for everyone during our pilot testing and subsequent user study. As our intention is to evaluate the effectiveness of the four techniques to highlight referents, we focus on the unidirectional brushing from the situated visualization to the referents.
@@ 4.4 Measures
Visual clutter. A highlighting technique ideally supports marking multiple referents without clutter, since scalability is often desired for both infovis [$*{}^{*}$22] and SitA [BPR20].
@@ 5.1 Task performance
For the overall linking completion time, which includes all tasks, we received 24 values per participant (4 techniques $\times$ 2 shelf layouts $\times$ 3 tasks). To aggregate these values for each technique and shelf layout per participant, we first averaged the values before calculating the overall mean and the respective confidence interval. The results of the overall linking completion time can be seen in 2(a). The Inside-FOV linking CT is generally faster than the Outside-FOV for all techniques. Color and Link perform similarly; Link is slightly better for the Inside-FOV, but worse for the Outside-FOV. While Outline resulted in similar results for the Inside-FOV, the results indicate a substantial performance drop for Outside-FOV targets. Outline is the only technique with a large gap between the two shelf layout conditions. Arrow is the worst for the Inside-FOV, and, second worst for Outside-FOV.
We also evaluated the linking completion time for the multi-selection task. We were specifically interested in this task, as multi-selection constitutes the "core discipline" of brushing and linking. The results are shown in 2(b). As before, the Inside-FOV has lower values than the Outside-FOV. Overall, Color performs best, closely followed by Link, as its Outside-FOV completion time is slightly higher. Again, Outline has a large gap between the shelf layout conditions, with the worst results for the Outside-FOV. Arrow is the worst for the Inside-FOV. By and large, the results for the multi-selection task are very similar to the overall results above, with a slightly stronger articulation of the main characteristics.
@@
We also measured errors as incorrectly selected products in single- or multi-selection tasks and removed duplicate errors caused by selecting the same incorrect product repeatedly. 3(a) shows the total count of errors among all participants. Color had the lowest number of errors and Arrow the most, with Outside-FOV being more prone to errors, especially for Link and Arrow. In addition, 3(b) shows for each error how far the correct product was in meters. From this result, we see that most of the errors made using Color and Outline were more than 1 m away, mainly due to incorrect brushing on the scatterplot. This was confirmed by reviewing the respective video footage. In contrast, for Link and Arrow, many errors were less than 0.5 m away, suggesting that participants actually mistook the product they needed to select—with some also due to incorrect brushing. Overall, Link and Arrow were more prone to errors when identifying a highlighted product.

### v3

@@ 1 Introduction
Situated analytics (SitA) involves the use of situated visualizations [WF09] to improve sense-making in physical contexts [ETM15, TWD18]. It is facilitated predominantly by the use of augmented reality (AR) [BKT22]. Willett et al. [WJD17] distinguish visualizations directly embedded with one particular physical referent from situated visualizations that indirectly link to a referent or the entire environment. Shin et al. [SBB23] note that most current SitA tools use the referent as a trigger to instantiate and modify visualizations. For instance, scanner applications augment real-world objects with additional information as they come into view. SitA is therefore reliant on the user’s viewpoint and interactions with the real world to determine what information to display.
Although this “physical world first” approach feels natural in the context of AR, it is inherently constrained in situations with many physical referents or referents that are spread far apart [LSS23]. Consider the classic supermarket scenario introduced by ElSayed et al. [ETM16], which features each product in the store having its own embedded visualization. While this scenario works well for products in view, it does not scale well to products outside of one’s view. Even with out-of-view labeling techniques [GLH18, LYBP23], the customer in one-to-one configuration would still need to look at and interpret the visualization of each individual product [LSS23]. The issue is exacerbated by the small field of view (FOV) of current state-of-the-art AR headsets.
@@ 2.1 Immersive and situated analytics
Immersive technologies, both VR and AR, are increasingly being used for immersive analytics [MSD18] of spatial data [FP21] and abstract data [KFS22]. Research has shown how the 3D environment can be used as an “immersive space to think” [LDG21], populating this space with multiple views and representations [RBR22]. However, care must be taken to consider the pitfalls of designing for immersive displays [ML14]. Designers of VR interfaces for analytics must take into account the characteristics of VR displays that differ from desktop computing, such as stereoscopic depth, angular resolution, field of view, or fatigue related to physical navigation and interaction [LKM17].
Recent work considers immersive analytics techniques with AR in physical environments, a new direction referred to as situated analytics (SitA) [ETM15, ETM16]. SitA is characterized by the ability to perform higher-level analytical reasoning aided by the user’s physical environment [SBB23], particularly when there is a semantic relationship between the data and its physical referents [WF09, WJD17]. While research has investigated how situated visualizations can be designed for SitA [LSS23, SBB23], considering human perception to leverage efficient visual processing—particularly in situated contexts—is a grand challenge [EBC21].
Designers of SitA systems (and AR interfaces in general) must handle the additional constraint that visual representations cannot always be positioned on a “clean slate” background. Kruijff et al. [KSF10] list several challenges related to visual perception in AR which are caused by the environment, image capture, display, and visual representation. Satkowski et al. [SD21] study how the real-world background affects perception in AR, and Assor et al. [APHD24] propose a design space of how to handle non-visible physical referents. Similarly, AR must always consider the interplay of visual representation with the existing and immutable appearance of the real world. Therefore, any visualization pipeline targeting AR needs to deal with the fusion of real and virtual visual attributes [ZLG21]. Addressing these challenges requires a wide range of visual encoding methods [SH16, chapter 7].
@@ 2.2 Brushing and linking
One of the core topics of this paper is brushing and linking [Rob07], which is the process of locating the corresponding items of interest in multiple views. Fundamentally, it comprises two sequential stages, wherein selections made in one visualization (brushing) are automatically shown in another visualization (linking) [Kei02]. Much research has focused on the brushing component, investigating interaction techniques to support the effective selection of data points in desktop computing [KPV18].
In immersive analytics, brushing and linking have seen (arguably limited) use, be it to compare multiple immersive coordinated views [AWG15, LPED20] or to support collaborative awareness [LHC21, SBDE23, SS22]. Most methods use simple 3D selection techniques (e.g., raycasting) or pure 2D selection techniques (e.g., touchscreen inputs). While we do not propose new brushing interaction techniques, we are the first (to the best of our knowledge) to explore brushing and linking in a SitA context—particularly when using physical referents as a coordinated view.
@@ 2.3 Attention guidance in augmented and virtual reality
In SitA, we assume that the linking component targets views and referents in the real world. Hence, it can be described as a visual search task for target objects in an AR display. Note that this work considers attention guidance as a means to support visual search and identification, and not as instructional cues for movement training [YLS24] or assembly tasks [PZB23]. Wolfe et al. [WVEG11] characterize visual search as the task of visually identifying a target among other, irrelevant objects (distractors). If there are too many distractors, a naive visual search becomes inefficient [Ver02]. We must therefore ensure that the visual search is tolerant of clutter and occlusion, since we cannot change the real world. There should be sufficient contrast and legibility between real and virtual objects. We have to be able to direct a user’s attention to the relevant objects, in particular, if these are out of view.
@@
According to Cockburn et al. [CKB09], conventional InfoVis methods to guide visual attention and search, such as overview + detail or focus + context, occupy a lot of screen space as they enlarge objects or introduce replicas of different sizes. As a major alternative, they mention cue-based approaches, which modify object rendering styles. Lin et al. [LYBP23] note that such cues are preferred in AR, since resizing or duplicating occupies too much of the natural view space. These cue-based techniques for highlighting focal objects can select from a variety of rendering attributes. A baseline is formed by highlights, which emphasize a target by modifying its color throughout the target area or as an outline [KPV18].
@@
In immersive displays, several more sophisticated techniques have been proposed to ensure that a given contrast is achieved [GAM18, KVZ13]. Contrast modulation, even below a detectable threshold [VMFS11], can pre-attentively guide the user. Other attributes that can be modified to highlight or attract attention include spatial frequencies (i.e., focus or blur) [KMH01], stereoscopy [KCWK20], or temporal frequencies [LSGB20].
@@
Cockburn et al. [CKB09] further mention proxies as cues. Arrow-shaped proxies are commonly used in AR [GLH18, TK06], most likely because they are simple to synthesize and easy to understand. Naturally, proxies are a popular means to guide toward out-of-view targets. In addition to arrows, various popular shapes include halos [BR03], funnels [BTOX06], compass-like [SL00] or radar-like widgets [JHPR11], and lines or trajectories [PLE19]. Previous work also compares the effect of various visual cues in VR and AR. Several authors [BSEN18, KKO14, TEM19, WGRF22] investigate the performance differences of visual search with respect to the field of view of the display. Other studies compare multiple attention guidance (i.e., highlighting) techniques with respect to the effectiveness and preservation of immersion [LSGB20], the effectiveness of multimodal cues [MTE20], or the influence of stationary and moving distractors [DARS23]. Moreover, our selection of techniques was influenced by the work of Whitlock et al. [WSS20] on the perception of fundamental graphical attributes in immersive analytics. However, none of these works investigate the use of AR or VR attention guidance in the context of brushing and linking. Therefore, our work seeks to bridge the gap between these two topics necessary to enable brushing and linking in SitA.
@@ 3.1 Scenario
SitA is relevant in many situations involving data and its physical context [BKT22]. Brushing and linking becomes useful when considering the actual appearance or location of physical referents. We therefore chose a supermarket scenario to situate our experimental prototype and the subsequent user study (Visual Highlighting for Situated Brushing and Linking). A supermarket is a popular scenario in the SitA literature [AWG15, ETM15, EST16]. Its defining feature is an abundance of grocery products (referents) laid out on shelves, which vary in layout, making the scenario well-suited for brushing and linking. We also considered similar environments, such as a library with books or a building façade with windows as referents [TOK16]. We built the supermarket using VR as a kind of information-rich virtual environment [BNC03, PKB05], serving as a simulated test environment that can be carefully controlled and replicated [RWBH09]. While using AR would be more ecologically valid, the use of VR avoids the technical challenges of tracking each product [CFSS23] or logistical challenges of renting or building a store. Previous work also suggested that findings in AR can be replicated in VR [GCOK21, LBHB09, LBBH10, LRM13], which further motivates and validates our use of a simulated supermarket in VR.
@@ 3.2 Highlighting techniques
Color is likely the de facto highlighting technique used in regular brushing and linking [KPV18, Rob07]. It mainly benefits from the pre-attentive “popout” effect, particularly when the new color highly contrasts with the background. Hence, we incorporate it into our prototype, changing the color of the product to a solid yellow when selected. While we had initially tried to use a semi-transparent yellow tint on the products instead, pilot tests showed that this mode was not able to sufficiently “pop-out” specific products: A red cereal box, when tinted yellow, simply looks like an orange cereal box, and not a red one that is highlighted.
@@
Outline is another common technique found in brushing and linking [GSL14, GGL14]. Moreover, it is also a popular technique used in visual highlighting to show that an object is “selected” [DMT18, SCZ20]. Although Color and Outline have similar characteristics in how they highlight targets, the actual appearance of a product is visible for Outline but not for Color. Sometimes a more subtle, non-obtrusive highlighting can be beneficial (e.g., for safety). We show outlines on the borders of the selected grocery products in camera space. We chose a width for the outline which, after pilot testing, we felt was sufficiently wide to see the outline, but thin enough not to obstruct neighboring products.
@@
Links are trajectories used to draw connections between entities [CC07]. In the context of brushing and linking, a link indicates that two connected data marks are the same underlying data record [KPV18]. The same premise holds in SitA, where a data mark on the situated visualization (i.e., the virtual tablet in Visual Highlighting for Situated Brushing and Linking) is explicitly connected to its associated referent. Particularly for the comparison and analysis across view, direct connections can facilitate tasks. To this end, we leverage a prior Unity implementation by Prouzeau et al. [PLE19] which draws visual links between pairs of objects in a 3D scene. We adjusted some of their script’s parameters to make the links more responsive to movement, because we allow the situated visualization to be moved by the user.
@@
Arrows are ubiquitous symbols used to direct attention. They are popular for attention guidance in AR, particularly towards objects or points of interest that are out of view [KR14, TK06, WGRF22, YLF20]. Thus, we consider how similar arrows could be used to highlight objects in SitA. For our study, we extended FlyingARrow [GLH18], which repeatedly instantiates 3D arrows in front of the user that then fly toward a target object. Static arrows like Wieland et al. [WGRF22] or Yu et al. [YLF20] spawn in the viewport, causing clutter when there are multiple targets and hence might be less beneficial in real-life scenarios than dynamic arrows [GLH18]. We also suspect that adding animations or motion makes targets more salient. Since Gruenefeld et al. [GLH18] claim a low workload, we suspected the dynamic arrow to be a reasonable choice for highlighting in SitA. Our adapted version of FlyingARrow is able to present multiple arrows concurrently and dynamically change the arrows to support on-the-fly brushing.
@@ 3.3 Brushing interactions
As the premise of our work involves both brushing and linking, our prototype needs faculties to enable brushing in a situated visualization. A virtual tablet attached to the non-dominant-hand controller (Visual Highlighting for Situated Brushing and Linking, center) presents a 2D scatterplot, generated with DXR [SLC19]. The $x$ and $y$ data dimensions of the scatterplot are determined by buttons on the right side of the tablet. The tablet’s size (40 cm by 31 cm) ensures that its contents are easily readable in VR, while still small enough to not be visually obtrusive.
We opt for a simple brushing approach, reminiscent of the brushing mechanism used in FIESTA [LHC21]. Brushing is performed by pointing the dominant-hand controller at the scatterplot and pressing either the index button to add to the selection or the grip button to remove it from the selection. We provide two modes for brushing: spherical and rectangular. The mode can be changed using the corresponding buttons on the tablet. Any brushed data marks on the scatterplot turn a solid yellow. To make selections even easier, we added simple range filters along the two data axes. These filters are represented as handles which can be dragged using raycast interactions, similar to ImAxes [CCD17]. Filtered data marks cannot be brushed, ensuring that no unwanted points are included. Step-by-step images of this situated brushing and linking process are shown in Figure 1. The tablet could be held comfortably at hip height without obscuring any referents, as was the case for everyone during our pilot testing and subsequent user study. As our intention is to evaluate the effectiveness of the four techniques to highlight referents, we focus on the unidirectional brushing from the situated visualization to the referents.
@@ 4.4 Measures
Visual clutter. A highlighting technique ideally supports marking multiple referents without clutter, since scalability is often desired for both infovis [RPA22] and SitA [BPR20].
@@ 5.1 Task performance
For the overall linking completion time, which includes all tasks, we received 24 values per participant (4 techniques $\times$ 2 shelf layouts $\times$ 3 tasks). To aggregate these values for each technique and shelf layout per participant, we first averaged the values before calculating the overall mean and the respective confidence interval. The results of the overall linking completion time can be seen in 3(a). The Inside-FOV linking CT is generally faster than the Outside-FOV for all techniques. Color and Link perform similarly; Link is slightly better for the Inside-FOV, but worse for the Outside-FOV. While Outline resulted in similar results for the Inside-FOV, the results indicate a substantial performance drop for Outside-FOV targets. Outline is the only technique with a large gap between the two shelf layout conditions. Arrow is the worst for the Inside-FOV, and, second worst for Outside-FOV.
We also evaluated the linking completion time for the multi-selection task. We were specifically interested in this task, as multi-selection constitutes the "core discipline" of brushing and linking. The results are shown in 3(b). As before, the Inside-FOV has lower values than the Outside-FOV. Overall, Color performs best, closely followed by Link, as its Outside-FOV completion time is slightly higher. Again, Outline has a large gap between the shelf layout conditions, with the worst results for the Outside-FOV. Arrow is the worst for the Inside-FOV. By and large, the results for the multi-selection task are very similar to the overall results above, with a slightly stronger articulation of the main characteristics.
@@
We also measured errors as incorrectly selected products in single- or multi-selection tasks and removed duplicate errors caused by selecting the same incorrect product repeatedly. 4(a) shows the total count of errors among all participants. Color had the lowest number of errors and Arrow the most, with Outside-FOV being more prone to errors, especially for Link and Arrow. In addition, 4(b) shows for each error how far the correct product was in meters. From this result, we see that most of the errors made using Color and Outline were more than 1 m away, mainly due to incorrect brushing on the scatterplot. This was confirmed by reviewing the respective video footage. In contrast, for Link and Arrow, many errors were less than 0.5 m away, suggesting that participants actually mistook the product they needed to select—with some also due to incorrect brushing. Overall, Link and Arrow were more prone to errors when identifying a highlighted product.